OpenStudy (anonymous):

Help ME !!!! Find < (theta) if cos(theta) =4/7 and sin is negative.

OpenStudy (anonymous):

I have to graph it.

OpenStudy (anonymous):

if cos(theta)=4/7, the value of theta is obtained by dividing cos on both sides so that you have theta=(4/7)/cos(this is like multiplying 4/7 by 1/cos. 1/cos is an inverted cosine By walking out using a mathematical table or a calculator, you will find the value of the inverted cosine of 4/7 to be 55.1501. Since the value of the sine of theta is a negative value, it means it is in the third and the fourth quadrants. Hence the values of theta as per the problem are 180+55.1501=235.1501 and 360-55.1501=304.8499 The values of theta are in degrees.

OpenStudy (anonymous):

how would it look if I graph it? @Klmusta

OpenStudy (ghostfire):

 this graph show sin is positive, cause sin(theta)=h/7, so you can either make h to negative or 7 to negative to get negative sin

OpenStudy (ghostfire):

but you can't make 7 to negative, so you have to make h to negative

OpenStudy (anonymous):

Finally if you make 7 to negative, then the change in y will be negative, i.e the gradient of the hypp will be negative. Remember that you're plotting the diagram on a cartesian plane.